OpenStudy (anonymous):

Use the three steps to solve the problem. Express the answers using decimals. It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? (Round answer to nearest tenth.) {Out___________________________________ hrs., back __________________________________ hrs.}

OpenStudy (anonymous):

i have a method, not sure if it is the shortest way, but it will get an answer

OpenStudy (anonymous):

let's call the distance between his house and his uncle's house \(D\) then since time is distance divided by rate, or \(T=\frac{D}{R}\) we know his time going is \[\frac{D}{20}\] and his time returning is \[\frac{D}{30}\] so his total time is \[\frac{D}{20}+\frac{D}{30}\] which you know is \(12\) so we can solve \[\frac{D}{20}+\frac{D}{30}=12\] for \(D\)

OpenStudy (anonymous):

Let t = time out...then 12 - t is the time back distance there = distance coming back speed there * time there = speed coming back * time coming back 20 t = 30 (12 - t).....solve for t

OpenStudy (anonymous):

guess my way was a bit cumbersome. go with @BangkokGarrett